A Bibliography of Papers in Lecture Notes in Computer Science (2003) (Part 1 of 4)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.06

Title word cross-reference

#11 [200].

7 [162]. GF (2) [73]. GF (p^{m}) [74]. GF(2) [173]. GF(2^{n}) [187]. μ [34, 40]. w [86].

-Chart-Based [34].

128-Bit [193].

ABC [6]. ABC/ADL [6]. Abstract [43, 22, 14]. Access [71]. Ada [29]. Address [198]. Address-Bit [198]. ADL [6]. AES [194, 175, 193]. Affine [95]. after [78]. against [197, 159, 198, 199, 86, 175, 179]. Agents [28]. Aggregate [118]. Agreement [127]. AI [110]. AKS [152]. Algebra [36, 29]. Algebraic [142, 113, 143, 135]. Algorithm [38, 145, 194, 176, 73, 185]. Algorithms [95, 116, 188, 179]. Analysis [197, 24, 198, 68, 58, 172, 40, 191, 119, 30]. Animation [21]. Anonymity [69]. Application [192, 185, 102, 175]. Applications [123, 74, 111, 89, 105]. Approach [15, 56, 18, 55, 156]. Arbitrarily [137]. Architecture [1, 12, 193]. Architectures [74, 11]. Arithmetic [74, 116, 176, 87]. Army [155]. Artifacts [181]. Assisted [88]. Assumption [79]. Assumptions [130, 97, 107, 138]. Asynchronous [180, 39]. Attack [192, 190, 115, 175, 191]. Attacking [189, 201]. Attacks [170, 142, 123, 134, 113, 143, 159, 199, 86, 172].

Authenticated [124, 111, 125, 83, 139, 156]. Authentication [161, 67]. Authenticity [163, 137]. Author [63]. Automata [45]. Automated [49]. Awareness [164].

B [49, 52]. Back [132]. Backdoors [91]. Based [106, 1, 130, 155, 161, 12, 76, 33, 181, 59, 125, 109, 34, 178, 201, 122, 27, 29, 9, 21, 23, 153, 71, 48, 7, 61]. Bases [135]. Basis [184]. Batch [24]. Better [190]. between [55, 127]. Bilinear [118, 70]. Birds [77]. Bit [198, 184, 193]. Block [178, 193, 75, 46]. Boolean [176]. Bound [157]. Bounded [136]. Bounds [158, 127]. Braid [145, 122]. Breaking [99]. Broadcast [81, 96]. Bubble [78]. Building [10, 16]. Bus [39]. Bytecode [36].

Cache [174]. Calculation [32]. Calculus [25, 35]. Camellia [193]. CAPTCHA [110]. Card [179]. CBC [66]. CCA2 [107]. CCA2-Secure [107]. Certificate [109]. Certificate-Based [109]. Challenges [138, 169]. Changes [47]. Changing [182]. Channel [166, 68, 199, 86, 170]. Characteristic [114]. Characterization [75]. Chart [34]. Chaumian [80]. Checker [41]. Checking [33, 45]. chip [39]. Chosen [165, 79]. Chosen-Ciphertext [165, 79]. Cipher [75]. Ciphers [113, 143, 178, 193]. Ciphertext [167, 165, 79]. Ciphertext-Only [167]. Circuit [181]. Circuits [180, 177, 159]. Circus} [62]. Class [187]. Closing [196]. Code [13, 161, 50]. Combiners [142]. commerce [12, 20]. Commitments [101]. Common [151]. Communication [167]. Compact [194]. Completeness [163, 61]. Complexity [184]. Component [13, 12, 10, 6]. Component-Based [12]. Components [8]. Composable [97, 147]. Composition [148, 6, 102]. Computable [136]. Computation [97, 129, 147, 96, 128, 98]. Computational [163]. Computers [174]. Computing [188, 4, 169]. Concealment [111]. Concept [27, 16]. Concept-Based [27]. Concurrent [178, 101, 35]. Conditional [59]. Confirmer [69]. Conjugacy [145]. Connections [112]. Constructing [136]. Construction [130, 107, 9]. Continuous [26]. Continuous-Time [26]. Control [71]. CORBA [7]. Correctness [32]. Countermeasure [198, 199]. Countermeasures [197]. Counting [114]. Critical [48]. CRL [40]. CRT [189]. Cryptanalysis [167, 144, 95, 135, 93, 171, 122, 174]. Cryptanalytic [168]. Cryptographic [74, 138]. Cryptography [150, 64, 78, 155, 76, 153, 71]. Cryptosystems [197, 135, 196, 136]. CSP [31, 14]. Cumulative [82]. Curve [197, 87, 196]. Curves [116, 114, 196].

Data [83]. Deadlock [29]. Decryption [140, 146]. Definition [49, 27]. Definitions [130]. DeKaRT [177]. Deniability [151]. Denotational [61]. Dependent [177]. Derandomization [150]. Description [59, 7]. Design [47, 56, 181, 195]. Detectable [96]. Detecting [29]. Detection [38, 178]. Determined [187]. Developing [41, 2]. Device [133]. Diagrams [52]. Differential [197, 198, 94, 175]. Diffie [145, 117, 156]. Digital [181, 72]. Directed [59]. Discrete [140]. Dishonest [128]. Distinction [16]. Distinguishers [94]. Distributed [56, 10, 81, 89]. Domain [13, 27, 58]. Domain-Specific [27]. Double [186]. Double-Size [186]. Doubling [190]. Downwards [190]. DPA [179]. Dynamic [33]. Dynamically [10].

E-commerce [12, 20]. ECPP [152]. Efficiency [128]. Efficient [74, 116, 129, 147, 185, 105, 154, 187, 101, 149, 195, 88].

Efficiently [141]. Elements [51]. Elliptic [116, 114, 87, 86, 196]. Embedding [15]. EMD [93]. Enabling [47]. Enciphering [160]. Encrypted [167, 118]. Encryption [106, 140, 108, 81, 147, 111, 65, 109, 163, 146, 122, 107, 80, 195, 179]. Endomorphisms [116]. Engineering [2]. Enigma [99]. Enough [162]. Enterprise [4]. Environment [182]. Equation [135]. Equations [173]. Equivalence [95]. Error [178]. Euclidean [186]. Evaluation [180, 87]. Event [37]. Evolutionary [28]. Exchange [124, 125, 139]. Experimental [172]. Exponentiation [187]. Exposure [134]. Extending [141]. Extracting [131]. Extraction [127]. Extractors [136].

Factoring [133]. Facts [99]. Failures [146]. Fair [98]. Fast [116, 143, 87, 86]. Faster [186, 168]. Fault [81, 175]. Feedback [113, 143]. Field [135]. Fields [187, 114]. Fighting [157]. Finite [187, 114]. First [172]. Fixed [112]. Flexible [199]. Flow [29]. FNLOG [23]. Formal [130, 37, 49, 24, 36, 8, 163, 4, 55, 2, 20, 44, 30, 28, 46]. Formalized [51]. Formalizing [53]. Formation [127]. Forward [64, 108]. Forward-Secure [108]. Forward-Security [64]. Foundations [130]. FPGA [194, 172]. Fractal [84]. Fraction [90]. Framework [125, 9]. Free [188]. Front [92]. Function [49]. Functions [157, 82].

Gap [196, 127]. Gauss [187]. GCD [188]. GCD-Free [188]. General [130, 107]. Generation [13, 91, 36, 89, 179, 82, 48]. Generator [183, 112, 181]. Generators [182]. Generic [16]. Geometric [83]. GGH [119]. GHS [115]. Given [90]. Goldreich [104]. Goubin [191]. GPS [192]. Grammar [13]. Graph [83]. Graphs [59]. Gröbner [135]. Group [130, 139, 131]. Groups [161, 122]. GSM [167]. GUI [37].

Handling [37]. Hard [110]. Hardware [47, 183, 173, 159, 19, 44, 193, 195, 75]. Hardware/Software [19]. Hashing [67]. Hellman [145, 117, 156]. HFE [135]. Hidden [171, 135]. Hierarchical [45]. History [158]. Homomorphic [147]. Hypercubic [119]. Hyperelliptic [197, 196].

Identity [155, 76]. Identity-Based [155, 76]. IKE [156]. Impact [78, 146]. Implementation [194, 181, 195]. Implemented [174]. Important [16]. Improved [116, 87]. Improvements [195]. Independence [158]. Index [63]. Instant [167]. Integrated [15]. Integrity [3]. Interaction [38]. Interception [166]. Interpretation [43]. Introducing [60]. Intrusion [65]. Intrusion-Resilient [65]. Inverses [188]. Invisibility [69]. Iteration [152].

Java [37]. Joint [148].

Kerckhoffs [79]. Key [106, 64, 123, 134, 108, 91, 124, 65, 125, 89, 177, 164, 163, 139, 66, 122, 107, 80, 127, 90, 100]. Key-Dependent [177]. KHAZAD [175]. Knife [155]. Knowledge [13, 103, 105, 101, 149, 100].

Language [42]. Languages [27]. Large [133]. Lattice [72, 149, 119]. Layered [1]. Left [185]. Length [80]. Length-Preserving [80]. Level [13, 179]. Levin [104]. Limitations [97]. Linear [95, 113, 143, 173, 94]. Locally [136]. Logarithms [140]. Logic [26, 179]. Logical [24]. Low [184]. Lower [158]. Luby [126, 162].

M2Z [42]. MAC [66, 156]. Machine

[50, 55]. Maintaining [3]. Majority [128]. Make [78]. Making [168]. malleable [105].

Many [126]. Many-Round [126]. Maps [118, 70, 82]. Markov [171]. Masking [176]. Matrix [161]. Matter [132, 92]. Matters [120]. Measurement [49]. Mechanization [15]. Mediated [76]. Memory [142, 157, 168, 86, 44]. Memory-Bound [157]. Merkle [84]. Message [161, 67]. Metamodeling [55]. Method [2, 199, 86]. Methodology [20]. Methods [4]. Micro [11]. Micro-architectures [11]. MIS [58]. MIST [90]. Mixes [80]. Mobile [25, 17]. Mode [160, 93]. Model [22, 10, 59, 8, 171, 21, 151, 57, 62, 136, 45, 100]. Model-Based [21]. Modeling [12, 51]. Modelling [37, 26]. Models [54, 53, 44]. Modular [186, 185, 188]. Montgomery [67]. Multi [170, 129, 96, 128]. Multi-channel [170]. Multi-party [129, 96, 128]. Multiparty [147]. Multiplication [186, 185]. Multiplications [86]. Multiplicative [154]. Multipliers [186, 184]. Multipurpose [155]. Mutations [48]. Myths [99].

NAF [86]. Natural [42]. Nearly [104]. Necessary [68]. Net [24]. Nets [7]. Network [25, 178]. Next [179]. Non [105, 61]. Non-malleable [105]. Non-redundancy [61]. NTRU [72, 146, 119]. NTRUSIGN [72]. Number [182, 183, 181]. Numbers [133].

Object [17, 55, 53, 14]. Object-Z [17, 55, 53, 14]. Objects [7]. Oblivious [141]. Obtaining [158]. Off [168]. On-chip [39]. One [152, 104, 154, 77]. One-Sided [104]. One-Way [154]. Only [167]. Ontology [58]. Ontology-Oriented [58]. Open [56]. Operation [93]. Operational [61]. Optimality [94]. Oracle [151]. Oriented [58]. Output [183]. Overview [17].

Padding [154]. Pairing [87, 71]. Paradigm [177]. Parallel [184, 5]. Parity [178]. Parity-Based [178]. Partial [134]. Partitioning [19]. Party [97, 98, 129, 96, 128]. Password [166, 124, 125]. Password-Authenticated [124]. Password-Based [125]. Pedersen [89]. ADL [6]. Glue [13]. PVS [15]. Software [19]. TLS [166, 201]. Performance [196, 75]. Periods [187]. Permutation [178, 154]. Permutations [126]. Petri [24, 7]. PKCS [200]. Plaintext [164, 105]. Points [49, 114]. Polynomial [106, 145, 102, 184]. Power [197, 198, 172, 191]. Power-Analysis [172]. Practical [140, 198]. Precondition [32]. Predicate [104]. Predict [183]. Predicting [112]. Preserving [80]. PRFs [123]. Primality [152]. Prime [67]. Privacy [137]. Private [64, 159, 90]. Private-Key [64]. Probing [159]. Problem [106, 145, 109, 117]. Problems [149, 110]. Process [29]. Programming [5]. Programs [22, 35]. Proofs [105, 149]. Protocol [89, 40, 102]. Protocols [103, 139, 156]. Provable [120]. Provably [124, 80, 88]. Prover [43, 32]. Provers [149]. Provides [86]. Proving [152]. Proxy [68]. PRPs [123]. Pseudo [126, 82]. Pseudo-Random [126, 82]. Public [106, 108, 65, 122, 107, 80, 100]. Public-Key [108, 65, 122, 107, 80, 100].

Quadratic [114]. Quality [2]. Quasi [114, 102]. Quasi-Polynomial [102].

Rackoff [126, 162]. RAISE [41]. Random [182, 183, 181, 126, 151, 82]. RDSA [121]. Real [22, 26, 35, 57, 30]. Real-Time [22, 26, 35, 57, 30]. Reasoning [44]. Reconfigurable [10, 195]. Reconstruction [106]. Reduction [185, 119]. redundancy [61]. Reference [151, 60]. Referential [3]. Refined [191]. Refinement [1, 36, 34, 16, 35, 60]. Registration [164].

Related [123]. Related-Key [123]. Relaxing [165]. Rendezvous [29]. Representation [24, 84]. Requirements [130, 59]. Research [58]. Resettable [100]. Resilient [65]. Results [172]. Rethinking [79]. Reuse [9]. Reversible [177]. Revisited [115]. Revocation [109]. Right [185]. Rijndael [195]. Rings [129]. RKA [123]. RKA-PRFs [123]. RKA-PRPs [123]. Round [152, 128, 126]. Rounds [162]. RSA [134, 91, 76, 189, 201, 77, 85, 90]. RSA-Based [201].

Safer [144]. Safety [50]. SAL [15]. SAL/ PVS [15]. SAM [30]. Scalable [139]. Scalar [86]. Scheduling [22]. Scheme [106, 70, 108, 117]. Schemes [131, 88]. Searching [83]. Secrecy [127]. Secret [127, 137]. Secret-Key [127]. Secure [182, 70, 108, 124, 89, 117, 68, 107, 80, 86, 98, 88]. Securing [159, 179]. Security [64, 165, 200, 121, 180, 146, 79, 126, 40, 162, 169, 120, 48, 110]. Security-Critical [48]. Seeing [90]. Semantic [57, 18]. Semantics [51, 60, 61]. Sequential [94]. Server [88]. Server-Assisted [88]. Service [38]. Sessions [201]. Set [97]. Set-up [97]. SFI [1]. Shallow [15]. Shared [82]. Shortcuts [85]. Shrinking [112]. Side [199, 86]. Sided [104]. SIGMA [156]. SIGn [156]. SIGn-and-MAc [156]. Signature [70, 117]. Signatures [130, 118, 69, 103, 72, 131, 68, 119, 88]. Signcryption [155, 77]. Simple [91, 76]. Simpler [107]. Simplified [130]. Simulatable [101]. Simulation [102]. Size [186]. Slicing [45]. Small [114, 86, 90]. Smart [179]. SOFL [2]. Software [1, 8, 2, 9, 44, 42, 28]. Solve [173]. SONET [46]. Soundness [61]. SPA [189]. Spam [157]. Sparse [173]. Specific [13, 27]. Specification [11, 13, 34, 52, 9, 21, 23, 39, 42, 14, 48, 28]. Specification-Based [9, 48].

Specifications [49, 16]. Specify [20].

Specifying [10]. Spell [41]. Spell-Checker [41]. SPN [175]. SSL [166, 201]. SSL/TLS [166, 201]. State [148, 55]. StateChart [52]. Statecharts [45]. STATEMATE [23]. Static [33, 51]. Statistical [149]. Stereotypes [99]. Stock [24]. Stone [77]. Storage [136]. Stream [113, 143]. Strengthening [103]. String [151]. Structure [29]. Structures [83, 175, 75]. Substitution [178]. Substitution-Permutation [178]. Support [32, 31]. Supporting [6]. Swiss [155]. Switching [176]. System [12, 46]. Systems [56, 10, 24, 173, 25, 26, 2, 54, 20, 48, 30].

Tajik [41]. TCOZ [33]. Technique

[23, 175]. Techniques [154]. Technology [179]. Telecom [46]. Temporal [26]. Test [48]. Testgraphs [21]. Tests [104]. Theorem [43, 32]. Theoretical [123]. Theories [5]. Three [11]. Three-Tiered [11]. Threshold [147, 124, 96, 88]. Tiered [11]. Time [22, 145, 114, 26, 168, 102, 35, 57, 62, 30]. Time-Memory [168]. Timing [192]. TMAC [66]. Tolerant [81]. Tool [31, 42]. Toolbox [95]. Torus [153]. Torus-Based [153]. Tracing [131]. Trade [168]. Trade-Off [168]. Tradeoffs [195]. Trading [24]. Traitor [131]. Transfers [141]. Transformation [55]. Transition [54]. Translating [42]. Trapdoor [154]. Traversal [84]. Treatment [123]. Tree [84]. True [182, 181]. Tweakable [160]. TWIRL [133]. Two [97, 13, 96, 66, 77, 98, 88]. Two-Key [66]. Two-Level [13]. Two-Party [97, 98]. Two-Threshold [96]. Type [192, 33, 50].

Ubiquitous [169]. UML

[56, 52, 51, 55, 54, 53, 31, 57, 45].

Unbalanced [189]. Unconditional [137].

Undeniable [69]. Unified [193]. Unify [54]. Unifying [5]. Unimodular [161]. Universal [148, 154]. Universally [97, 147]. Upper [158]. Upwards [190]. Use [156, 16]. Using [13, 116, 41, 135, 189, 103, 72, 2, 54, 77, 21, 199, 71, 48, 110].

Validation [23]. Verifiable [140]. Verifiably [118]. Verification [47, 46]. Verilog [19, 61]. Very [194]. via [152, 164, 60]. Visualization [33]. Visualizing [31].

Way [154]. Weak [163, 137, 100]. Web [18, 3]. Weil [87]. Width [86]. Width- [86]. Window [199]. without [97]. Workflows [24]. World [78]. Wrapper [13]. Wrapper/ Glue [13].

XL [73]. XML [33]. XML-Based [33].

Z [18, 55, 53, 42, 14, 17]. Zero [103, 101, 149, 100]. Zero-Knowledge [103, 101, 149, 100].

References

Back:2003:SRB

[1] Ralph-Johan Back. SFI: A refinement based layered software architecture. Lecture Notes in Computer Science, 2495:1–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950001.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950001.pdf.

Liu:2003:DQS

[2] Shaoying Liu. Developing quality software systems using the SOFL formal engineering method. Lecture Notes in Computer Science, 2495:3– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950003.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950003.pdf.

Hale:2003:MRI

[3] Mark A. Hale. Maintaining referential integrity on the Web. Lecture Notes in Computer Science, 2495:20– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950020.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950020.pdf.

Jullig:2003:FME

[4] Richard Jüllig. Formal methods in enterprise computing. Lecture Notes in Computer Science, 2495:22– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950022.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950022.pdf.

Woodcock:2003:UTP

[5] Jim Woodcock and Arthur Hughes. Unifying theories of parallel programming. Lecture Notes in Computer Science, 2495:24–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950024.htm;   http://link.

springer.de/link/service/series/
0558/papers/2495/24950024.pdf.

Mei:2003:AAA

[6] Hong Mei, Feng Chen, Qianxiang Wang, and Yao-Dong Feng. ABC/ ADL: An ADL supporting component composition. Lecture Notes in Computer Science, 2495:38–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950038.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950038.pdf.

Zheng:2003:DCO

[7] Hong Zheng and Shi xian Li. The description of CORBA objects based on Petri nets. Lecture Notes in Computer Science, 2495:48–??, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2495/24950048.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2495/24950048.pdf.

Heisel:2003:TFM

[8] Maritta Heisel, Thomas Santen, and Jeanine Souquières. Toward a formal model of software components. Lecture Notes in Computer Science, 2495:57–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950057.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950057.pdf.

Liu:2003:SBS

[9] Jing Liu, Huaikou Miao, and Xiaolei Gao. A specification-based software construction framework for reuse. Lecture Notes in Computer Science, 2495:69–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950069.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950069.pdf.

Chen:2003:SCM

[10] Xuejun Chen. Specifying a component model for building dynamically reconfigurable distributed systems. Lecture Notes in Computer Science, 2495:80– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950080.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950080.pdf.

Alagar:2003:TTS

[11] Vasu Alagar and Ralf Lämmel. Threetiered specification of micro-architectures. Lecture Notes in Computer Science, 2495:92–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950092.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950092.pdf.

Chang:2003:MAC

[12] Jiayue Chang and Huadong Ma. Modeling the architecture for componentbased E-commerce system. Lecture Notes in Computer Science, 2495:98– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950098.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950098.pdf.

Cao:2003:CSW

[13] Fei Cao, Barrett R. Bryant, Rajeev R. Raje, Mikhail Auguston, Andrew M. Olson, and Carol C. Burt. Component specification and wrapper/glue code generation with two-level grammar using domain specific knowledge. Lecture Notes in Computer Science, 2495:103–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950103.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950103.pdf.

Smith:2003:ASO

[14] Graeme Smith and John Derrick. Abstract specification in Object-Z and CSP. Lecture Notes in Computer Science, 2495:108–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2495/24950108.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2495/24950108.pdf.

Attiogbe:2003:MIA

[15] J. Christian Attiogbé. Mechanization of an integrated approach: Shallow embedding into SAL/PVS. Lecture Notes in Computer Science, 2495:

120–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950120.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950120.pdf.

Musser:2003:CUC

[16] David R. Musser and Zhiqing Shao. Concept use or concept refinement: An important distinction in building generic specifications. Lecture Notes in Computer Science, 2495:132– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950132.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950132.pdf.

Taguchi:2003:OMO

[17] Kenji Taguchi and Jin Song Dong. An overview of mobile object-Z. Lecture Notes in Computer Science, 2495: 144–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950144.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950144.pdf.

Dong:2003:ZAS

[18] Jin Song Dong, Jing Sun, and Hai Wang. Z approach to Semantic Web. Lecture Notes in Computer Science, 2495:156–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950156.htm;   http://link.

springer.de/link/service/series/
0558/papers/2495/24950156.pdf.

Qin:2003:HSP

[19] Shengchao Qin, Jifeng He, Zongyan Qiu, and Naixiao Zhang. Hardware/ software partitioning in Verilog. Lecture Notes in Computer Science, 2495: 168–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950168.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950168.pdf.

Pereira:2003:FMS

[20] Adriano Pereira, Mark Song, Gustavo Gorgulho, Wagner Meira Jr., and Sérgio Campos. A formal methodology to specify E-commerce systems. Lecture Notes in Computer Science, 2495:180–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950180.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950180.pdf.

Miller:2003:MBS

[21] Tim Miller and Paul Strooper. Modelbased specification animation using testgraphs. Lecture Notes in Computer Science, 2495:192–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2495/24950192.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2495/24950192.pdf.

Arenas:2003:AMS

[22] Alvaro E. Arenas. An abstract model for scheduling real-time programs. Lecture Notes in Computer Science, 2495: 204–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950204.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950204.pdf.

Mosbahi:2003:SVT

[23] Olfa Mosbahi, Leila Jemni, Samir Ben Ahmed, and Jacques Jaray. A specification and validation technique based on STATEMATE and FNLOG. Lecture Notes in Computer Science, 2495: 216–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950216.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950216.pdf.

Du:2003:FRA

[24] Yuyue Du and Changjun Jiang. Formal representation and analysis of batch stock trading systems by logical Petri net workflows. Lecture Notes in Computer Science, 2495:221– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950221.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950221.pdf.

Huang:2003:CMN

[25] Jinfeng Huang, Ad Verschueren, Henri Aalderink, and Johan Lukkien. A calculus for mobile network systems. Lecture Notes in Computer Science, 2495:226–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950226.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950226.pdf.

Li:2003:MRT

[26] Guangyuan Li and Zhisong Tang. Modelling real-time systems with continuoustime temporal logic. Lecture Notes in Computer Science, 2495:231–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950231.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950231.pdf.

Liu:2003:CBD

[27] Ying Liu and Naixiao Zhang. On concept-based definition of domainspecific languages. Lecture Notes in Computer Science, 2495:237–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950237.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950237.pdf.

Zhu:2003:FSE

[28] Hong Zhu. Formal specification of evolutionary software agents. Lecture Notes in Computer Science, 2495: 249–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.

                                        de/link/service/series/0558/bibs/
                                        2495/24950249.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950249.pdf.

Liu:2003:DDA

[29] Yuan Liu, Baowen Xu, and Zhenqiang Chen. Detecting deadlock in Ada rendezvous flow structure based on process algebra. Lecture Notes in Computer Science, 2495:262–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950262.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950262.pdf.

Yu:2003:FAR

[30] Huiqun Yu, Xudong He, Yi Deng, and Lian Mo. Formal analysis of real-time systems with SAM. Lecture Notes in Computer Science, 2495: 275–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950275.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950275.pdf.

Ng:2003:TSV

[31] Muan Yong Ng and Michael Butler. Tool support for visualizing CSP in UML. Lecture Notes in Computer Science, 2495:287–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950287.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950287.pdf.

Celiku:2003:TPS

[32] Orieta Celiku and Joakim von Wright. Theorem prover support for precondition and correctness calculation. Lecture Notes in Computer Science, 2495: 299–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950299.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950299.pdf.

Dong:2003:XBS

[33] Jin Song Dong, Yuan Fang Li, Jing Sun, Jun Sun, and Hai Wang. XMLbased static type checking and dynamic visualization for TCOZ. Lecture Notes in Computer Science, 2495: 311–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950311.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950311.pdf.

Goldson:2003:CBS

[34] Doug Goldson, Greg Reeve, and Steve Reeves. μ-chart-based specification and refinement. Lecture Notes in Computer Science, 2495:323– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950323.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950323.pdf.

Peuker:2003:TRC

[35] Sibylle Peuker and Ian Hayes. Towards a refinement calculus for concurrent real-time programs. Lecture Notes in Computer Science, 2495:335– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950335.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950335.pdf.

Duran:2003:RAF

[36] Adolfo Duran, Ana Cavalcanti, and Augusto Sampaio. Refinement algebra for formal bytecode generation. Lecture Notes in Computer Science, 2495:347–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950347.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950347.pdf.

Chen:2003:FMJ

[37] Jessica Chen. Formal modelling of Java GUI event handling. Lecture Notes in Computer Science, 2495:359– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950359.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950359.pdf.

Cavalli:2003:NAS

[38] Ana Cavalli and Stéphane Maag. A new algorithm for service interaction detection. Lecture Notes in Computer Science, 2495:371–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.

de/link/service/series/0558/bibs/
2495/24950371.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950371.pdf.

Plosila:2003:SAC

[39] Juha Plosila and Tiberiu Seceleanu. Specification of an asynchronous on-chip bus. Lecture Notes in Computer Science, 2495:383–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950383.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950383.pdf.

Pang:2003:ASP

[40] Jun Pang. Analysis of a security protocol in μCRL. Lecture Notes in Computer Science, 2495:396–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950396.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950396.pdf.

Davrondjon:2003:DSC

[41] Gafurov Davrondjon and Tomasz Janowski. Developing a spell-checker for Tajik using RAISE. Lecture Notes in Computer Science, 2495:401– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950401.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950401.pdf.

Shukur:2003:MTT

[42] Zarina Shukur, Abdullah Md. Zin, and Ainita Ban. M2Z: A tool for translating a natural language software specification into Z. Lecture Notes in Computer Science, 2495:406– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950406.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950406.pdf.

Anderson:2003:AIT

[43] Hugh Anderson. Abstract interpretation with a theorem prover. Lecture Notes in Computer Science, 2495: 411–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950411.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950411.pdf.

Roychoudhury:2003:FRA

[44] Abhik Roychoudhury. Formal reasoning about hardware and software memory models. Lecture Notes in Computer Science, 2495:423–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950423.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950423.pdf.

Wang:2003:SHA

[45] Ji Wang, Wei Dong, and Zhi-Chang Qi. Slicing hierarchical automata for model checking UML statecharts. Lecture Notes in Computer Science, 2495: 435–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950435.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950435.pdf.

Zobair:2003:FVS

[46] M. Hasan Zobair and Sofiène Tahar. Formal verification of a SONET telecom system block. Lecture Notes in Computer Science, 2495:447–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950447.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950447.pdf.

Abdel-Hamid:2003:EHV

[47] Amr T. Abdel-Hamid, Sofiène Tahar, and John Harrison. Enabling hardware verification through design changes. Lecture Notes in Computer Science, 2495:459–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950459.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950459.pdf.

Wimmel:2003:SBT

[48] Guido Wimmel and Jan Jürjens. Specification-based test generation for security-critical systems using mutations. Lecture Notes in Computer Science, 2495:471–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2495/24950471.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2495/24950471.pdf.

Diab:2003:FDF

[49] Hassan Diab, Marc Frappier, and Richard St-Denis. A formal definition of function points for automated measurement of B specifications. Lecture Notes in Computer Science, 2495: 483–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950483.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950483.pdf.

Guo:2003:MCT

[50] Fan Guo, YiYun Chen, and Rong-Gui Hu. Machine code type safety. Lecture Notes in Computer Science, 2495:495–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950495.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950495.pdf.

Jiang:2003:FSS

[51] Yan-Bing Jiang, Wei-Zhong Shao, Zhi-Yi Ma, and Yao-Dong Feng. On the formalized semantics of static modeling elements in UML. Lecture Notes in Computer Science, 2495:500– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/

2495/24950500.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950500.pdf.

Hammad:2003:BSU

[52] Ahmed Hammad, Bruno Tatibouët, Jean-Christophe Voisinet, and Weiping Wu. From a B specification to UML StateChart diagrams. Lecture Notes in Computer Science, 2495:511– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950511.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950511.pdf.

Miao:2003:FUM

[53] Huaikou Miao, Ling Liu, and Li Li. Formalizing UML models with Object-Z. Lecture Notes in Computer Science, 2495:523–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950523.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950523.pdf.

Liu:2003:UTS

[54] Zhiming Liu, Xiaoshan Li, and Jifeng He. Using transition systems to unify UML models. Lecture Notes in Computer Science, 2495:535–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950535.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950535.pdf.

Kim:2003:FMA

[55] Soon-Kyeong Kim and David Carrington. A formal metamodeling approach to a transformation between the UML state machine and Object-Z. Lecture Notes in Computer Science, 2495:548–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950548.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950548.pdf.

Bordbar:2003:UAD

[56] Behzad Bordbar, John Derrick, and Gill Waters. A UML approach to the design of open distributed systems. Lecture Notes in Computer Science, 2495:561–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950561.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950561.pdf.

Shankar:2003:SMR

[57] Subash Shankar. A semantic model of real-time UML. Lecture Notes in Computer Science, 2495:573–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950573.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950573.pdf.

Ming:2003:ROO

[58] Zhong Ming, Shi xian Li, and Xiu rong Fang. Research on ontologyoriented domain analysis on MIS. Lecture Notes in Computer Science, 2495: 578–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950578.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950578.pdf.

Gan:2003:RDM

[59] Zaobin Gan, Chuanbo Chen, and Xiandeng Pei. A requirements description model based on conditional directed graphs. Lecture Notes in Computer Science, 2495:583–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950583.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950583.pdf.

Smith:2003:IRS

[60] Graeme Smith. Introducing reference semantics via refinement. Lecture Notes in Computer Science, 2495: 588–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2495/24950588.htm;   http://link.
springer.de/link/service/series/
0558/papers/2495/24950588.pdf.

Zhu:2003:SCN

[61] Huibiao Zhu, Jonathan P. Bowen, and Jifeng He. Soundness, completeness and non-redundancy of operational semantics for Verilog based on denotational semantics. Lecture Notes in Computer Science, 2495:600– ??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950600.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950600.pdf.

Sherif:2003:TTM

[62] Adnan Sherif and Jifeng He. Towards a time model for Circus. Lecture Notes in Computer Science, 2495: 613–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2495/24950613.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2495/24950613.pdf.

Anonymous:2003:AIg

[63] Anonymous. Author index. Lecture Notes in Computer Science, 2495: 625–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2495/2495auth.pdf.

Bellare:2003:FSP

[64] Mihir Bellare and Bennet Yee. Forwardsecurity in private-key cryptography. Lecture Notes in Computer Science, 2612:1–18, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dodis:2003:IRP

[65] Yevgeniy Dodis, Matt Franklin, Jonathan Katz, Atsuko Miyaji, and Moti Yung. Intrusion-resilient public-key encryption. Lecture Notes in Computer Science, 2612:19–32, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:2003:TTK

[66] Kaoru Kurosawa and Tetsu Iwata. TMAC: Two-key CBC MAC. Lecture Notes in Computer Science, 2612:33–49, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whiting:2003:MPH

[67] Douglas L. Whiting and Michael J. Sabin. Montgomery prime hashing for message authentication. Lecture Notes in Computer Science, 2612:50–67, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:2003:APS

[68] Jung-Yeun Lee, Jung Hee Cheon, and Seungjoo Kim. An analysis of proxy signatures: Is a secure channel necessary? Lecture Notes in Computer Science, 2612:68–79, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galbraith:2003:IAU

[69] Steven D. Galbraith and Wenbo Mao. Invisibility and anonymity of undeniable and confirmer signatures. Lecture Notes in Computer Science, 2612:80–97, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boneh:2003:SSS

[70] Dan Boneh, Ilya Mironov, and Victor Shoup. A secure signature scheme from bilinear maps. Lecture Notes in Computer Science, 2612:98–110, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smart:2003:ACU

[71] Nigel P. Smart. Access control using pairing based cryptography. Lecture Notes in Computer Science, 2612: 111–121, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffstein:2003:NDS

[72] Jeffrey Hoffstein, Nick Howgrave-Graham, Jill Pipher, Joseph H. Silverman, and William Whyte. NTRUSIGN: Digital signatures using the NTRU lattice. Lecture Notes in Computer Science, 2612:122–140, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courtois:2003:AXA

[73] Nicolas T. Courtois and Jacques Patarin. About the XL algorithm over GF (2). Lecture Notes in Computer Science, 2612:141–157, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertoni:2003:EAA

[74] Guido Bertoni, Jorge Guajardo, Sandeep Kumar, Gerardo Orlando, Christof Paar, and Thomas Wollinger. Efficient GF (p^{m}) arithmetic architectures for cryptographic applications. Lecture Notes in Computer Science, 2612: 158–175, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xiao:2003:HPC

[75] Lu Xiao and Howard M. Heys. Hardware performance characterization of block cipher structures. Lecture Notes in Computer Science, 2612:176–192, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:2003:SIB

[76] Xuhua Ding and Gene Tsudik. Simple identity-based cryptography with mediated RSA. Lecture Notes in Computer Science, 2612:193–210, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malone-Lee:2003:TBO

[77] John Malone-Lee and Wenbo Mao. Two birds one stone: Signcryption using RSA. Lecture Notes in Computer Science, 2612:211–225, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berson:2003:CAB

[78] Tom Berson. Cryptography after the bubble: How to make an impact on the world. Lecture Notes in Computer Science, 2612:226, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:2003:RCC

[79] Seungjoo Kim, Masahiro Mambo, and Yuliang Zheng. Rethinking chosenciphertext security under Kerckhoffs’ assumption. Lecture Notes in Computer Science, 2612:227–243, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:2003:PSP

[80] Bodo Möller. Provably secure public-key encryption for length-preserving Chaumian mixes. Lecture Notes in Computer Science, 2612:244–262, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DArco:2003:FTD

[81] Paolo D’Arco and Douglas R. Stinson. Fault tolerant and distributed broadcast encryption. Lecture Notes in Computer Science, 2612:263–280, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:2003:SGP

[82] Huaxiong Wang and Josef Pieprzyk. Shared generation of pseudo-random functions with cumulative maps. Lecture Notes in Computer Science, 2612: 281–294, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goodrich:2003:ADS

[83] Michael T. Goodrich, Roberto Tamassia, Nikos Triandopoulos, and Robert Cohen. Authenticated data structures for graph and geometric searching. Lecture Notes in Computer Science, 2612: 295–313, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobsson:2003:FMT

[84] Markus Jakobsson, Tom Leighton, Silvio Micali, and Michael Szydlo. Fractal Merkle tree representation and traversal. Lecture Notes in Computer Science, 2612:314–326, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shamir:2003:RS

[85] Adi Shamir. RSA shortcuts. Lecture Notes in Computer Science, 2612:327, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okeya:2003:WNM

[86] Katsuyuki Okeya and Tsuyoshi Takagi. The width-w NAF method provides small memory and fast elliptic scalar multiplications secure against side channel attacks. Lecture Notes in Computer Science, 2612:328–342, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montgomery:2003:FEC

[87] Peter L. Montgomery, Kirsten Eisenträger, and Kristin Lauter. Fast elliptic curve arithmetic and improved Weil pairing evaluation. Lecture Notes in Computer Science, 2612:343–354, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:2003:TEP

[88] Shouhuai Xu and Ravi Sandhu. Two efficient and provably secure schemes for server-assisted threshold signatures. Lecture Notes in Computer Science, 2612:355–372, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gennaro:2003:SAP

[89] Rosario Gennaro, Stanislaw Jarecki, Hugo Krawczyk, and Tal Rabin. Secure applications of Pedersen’s distributed key generation protocol. Lecture Notes in Computer Science, 2612:373–390, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walter:2003:STM

[90] Colin D. Walter. Seeing through MIST given a small fraction of an RSA private key. Lecture Notes in Computer Science, 2612:391–402, 2003. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Crepeau:2003:SBR

[91] Claude Crépeau and Alain Slakmon. Simple backdoors for RSA key generation. Lecture Notes in Computer Science, 2612:403–416, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:2003:FM

[92] Anonymous. Front matter. Lecture Notes in Computer Science, 2656: i–xiv, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/bfm:978-
                                        3-540-39200-2/1.pdf.

Joux:2003:CEM

[93] Antoine Joux. Cryptanalysis of the EMD mode of operation. Lecture Notes in Computer Science, 2656:1–16, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/3-540-39200-9_1.pdf.

Junod:2003:OLD

[94] Pascal Junod. On the optimality of linear, differential, and sequential distinguishers. Lecture Notes in Computer Science, 2656:17–32, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_2.pdf.

Biryukov:2003:TCL

[95] Alex Biryukov, Christophe De Cannière, An Braeken, and Bart Preneel. A toolbox for cryptanalysis: Linear and affine equivalence algorithms. Lecture Notes in Computer Science, 2656:33–50, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/3-540-39200-9_3.pdf.

Fitzi:2003:TTB

[96] Matthias Fitzi, Martin Hirt, Thomas Holenstein, and Jürg Wullschleger. Twothreshold broadcast and detectable multi-party computation. Lecture Notes in Computer Science, 2656:51–67, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/3-540-39200-9_4.pdf.

Canetti:2003:LUC

[97] Ran Canetti, Eyal Kushilevitz, and Yehuda Lindell. On the limitations of universally composable two-party computation without set-up assumptions. Lecture Notes in Computer Science, 2656:68–86, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_5.pdf.

Pinkas:2003:FST

[98] Benny Pinkas. Fair secure two-party computation. Lecture Notes in Computer Science, 2656:87–105, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/3-540-39200-9_6.pdf.

Gaj:2003:FME

[99] Kris Gaj and Arkadiusz Orłowski. Facts and myths of Enigma: Breaking stereotypes. Lecture Notes in Computer Science, 2656:106–122, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_7.pdf.

Zhao:2003:RZK

[100] Yunlei Zhao, Xiaotie Deng, C. H. Lee, and Hong Zhu. Resettable zero-knowledge in the weak public-key model. Lecture Notes in Computer Science, 2656:123–139, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_8.pdf.

Micciancio:2003:SCE

[101] Daniele Micciancio and Erez Petrank. Simulatable commitments and efficient concurrent zero-knowledge. Lecture Notes in Computer Science, 2656: 140–159, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_9.pdf.

Pass:2003:SQP

[102] Rafael Pass. Simulation in quasipolynomial time, and its application to protocol composition. Lecture Notes in Computer Science, 2656:160–176, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/3-540-39200-9_10.pdf.

Garay:2003:SZK

[103] Juan A. Garay, Philip MacKenzie, and Ke Yang. Strengthening zeroknowledge protocols using signatures. Lecture Notes in Computer Science, 2656:177–194, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_11.pdf.

Hast:2003:NOS

[104] Gustav Hast. Nearly one-sided tests and the Goldreich–Levin predicate. Lecture Notes in Computer Science, 2656: 195–210, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_12.pdf.

Katz:2003:ENM

[105] Jonathan Katz. Efficient and nonmalleable proofs of plaintext knowledge and applications. Lecture Notes in Computer Science, 2656:211–228, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/3-540-39200-9_13.pdf.

Augot:2003:PKE

[106] Daniel Augot and Matthieu Finiasz.

A public key encryption scheme based on the polynomial reconstruction problem. Lecture Notes in Computer Science, 2656:229–240, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_14.pdf.

Lindell:2003:SCC

[107] Yehuda Lindell. A simpler construction of CCA2-secure public-key encryption under general assumptions. Lecture Notes in Computer Science, 2656: 241–254, 2003. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_15.pdf.

Canetti:2003:FSP

[108] Ran Canetti, Shai Halevi, and Jonathan Katz. A forward-secure public-key encryption scheme. Lecture Notes in Computer Science, 2656:255–271, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/3-540-39200-9_16.pdf.

Gentry:2003:CBE

[109] Craig Gentry. Certificate-based encryption and the certificate revocation problem. Lecture Notes in Computer Science, 2656:272–293, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_17.pdf.

vonAhn:2003:CUH

[110] Luis von Ahn, Manuel Blum, Nicholas J. Hopper, and John Langford. CAPTCHA: Using hard AI problems for security. Lecture Notes in Computer Science, 2656:294–311, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_18.pdf.

Dodis:2003:CAA

[111] Yevgeniy Dodis and Jee Hea An. Concealment and its applications to authenticated encryption. Lecture Notes in Computer Science, 2656:312–329, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/3-540-39200-9_19.pdf.

Ekdahl:2003:PSG

[112] Patrik Ekdahl, Willi Meier, and Thomas Johansson. Predicting the shrinking generator with fixed connections. Lecture Notes in Computer Science, 2656:330–344, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_20.pdf.

Courtois:2003:AAS

[113] Nicolas T. Courtois and Willi Meier. Algebraic attacks on stream ciphers with linear feedback. Lecture Notes in Computer Science, 2656:345–359, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/3-540-39200-9_21.pdf.

Lercier:2003:CPE

[114] Reynald Lercier and David Lubicz.

Counting points on elliptic curves over finite fields of small characteristic in quasi quadratic time. Lecture Notes in Computer Science, 2656:360–373, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/3-540-39200-9_22.pdf.

Hess:2003:GAR

[115] Florian Hess. The GHS attack revisited. Lecture Notes in Computer Science, 2656:374–387, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_23.pdf.

Ciet:2003:IAE

[116] Mathieu Ciet, Tanja Lange, Francesco Sica, and Jean-Jacques Quisquater. Improved algorithms for efficient arithmetic on elliptic curves using fast endomorphisms. Lecture Notes in Computer Science, 2656:388–400, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/3-540-39200-9_24.pdf.

Goh:2003:SSS

[117] Eu-Jin Goh and Stanisław Jarecki. A signature scheme as secure as the Diffie– Hellman problem. Lecture Notes in Computer Science, 2656:401–415, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/3-540-39200-9_25.pdf.

Boneh:2003:AVE

[118] Dan Boneh, Craig Gentry, Ben Lynn, and Hovav Shacham. Aggregate and verifiably encrypted signatures from bilinear maps. Lecture Notes in Computer Science, 2656:416–432, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_26.pdf.

Szydlo:2003:HLR

[119] Michael Szydlo. Hypercubic lattice reduction and analysis of GGH and NTRU signatures. Lecture Notes in Computer Science, 2656:433–448, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        3-540-39200-9_27.pdf.

Stern:2003:WPS

[120] Jacques Stern. Why provable security matters? Lecture Notes in Computer Science, 2656:449–461, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_28.pdf.

Fouque:2003:SR

[121] Pierre-Alain Fouque and Guillaume Poupard. On the security of RDSA. Lecture Notes in Computer Science, 2656:462–476, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_29.pdf.

Lee:2003:CPK

[122] Eonkyung Lee and Je Hong Park. Cryptanalysis of the public-key encryption based on braid groups. Lecture Notes in Computer Science, 2656:477–490, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/3-540-39200-9_30.pdf.

Bellare:2003:TTR

[123] Mihir Bellare and Tadayoshi Kohno. A theoretical treatment of related-key attacks: RKA-PRPs, RKA-PRFs, and applications. Lecture Notes in Computer Science, 2656:491–506, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_31.pdf.

DiRaimondo:2003:PST

[124] Mario Di Raimondo and Rosario Gennaro. Provably secure threshold password-authenticated key exchange. Lecture Notes in Computer Science, 2656:507–523, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_32.pdf.

Gennaro:2003:FPB

[125] Rosario Gennaro and Yehuda Lindell. A framework for password-based authenticated key exchange. Lecture Notes in Computer Science, 2656:524–543, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/3-540-39200-9_33.pdf.

Maurer:2003:SMR

[126] Ueli Maurer and Krzysztof Pietrzak. The security of many-round Luby– Rackoff pseudo-random permutations. Lecture Notes in Computer Science, 2656:544–561, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_34.pdf.

Renner:2003:NBS

[127] Renato Renner and Stefan Wolf. New bounds in secret-key agreement: The gap between formation and secrecy extraction. Lecture Notes in Computer Science, 2656:562–577, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_35.pdf.

Katz:2003:REM

[128] Jonathan Katz, Rafail Ostrovsky, and Adam Smith. Round efficiency of multiparty computation with a dishonest majority. Lecture Notes in Computer Science, 2656:578–595, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_36.pdf.

Cramer:2003:EMP

[129] Ronald Cramer, Serge Fehr, Yuval Ishai, and Eyal Kushilevitz. Efficient multi-party computation over rings. Lecture Notes in Computer Science, 2656:596–613, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_37.pdf.

Bellare:2003:FGS

[130] Mihir Bellare, Daniele Micciancio, and Bogdan Warinschi. Foundations of group signatures: Formal definitions, simplified requirements, and a construction based on general assumptions. Lecture Notes in Computer Science, 2656:614–629, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
3-540-39200-9_38.pdf.

Kiayias:2003:EGS

[131] Aggelos Kiayias and Moti Yung. Extracting group signatures from traitor tracing schemes. Lecture Notes in Computer Science, 2656:630–648, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/3-540-39200-9_39.pdf.

Anonymous:2003:BM

[132] Anonymous. Back matter. Lecture Notes in Computer Science, 2656:649, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/bbm:978-3-540-39200-2/1.pdf.

Shamir:2003:FLN

[133] Adi Shamir and Eran Tromer. Factoring large numbers with the TWIRL device. Lecture Notes in Computer Science, 2729:1–26, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blomer:2003:NPK

[134] Johannes Blömer and Alexander May. New partial key exposure attacks on RSA. Lecture Notes in Computer Science, 2729:27–43, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faugere:2003:ACH

[135] Jean-Charles Faugère and Antoine Joux. Algebraic cryptanalysis of Hidden Field Equation (HFE) cryptosystems using Gröbner bases. Lecture Notes in Computer Science, 2729:44–60, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vadhan:2003:CLC

[136] Salil P. Vadhan. On constructing locally computable extractors and cryptosystems in the bounded storage model. Lecture Notes in Computer Science, 2729:61–77, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Renner:2003:UAP

[137] Renato Renner and Stefan Wolf. Unconditional authenticity and privacy from an arbitrarily weak secret. Lecture Notes in Computer Science, 2729:78–95, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naor:2003:CAC

[138] Moni Naor. On cryptographic assumptions and challenges. Lecture Notes in Computer Science, 2729:96–109, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katz:2003:SPA

[139] Jonathan Katz and Moti Yung. Scalable protocols for authenticated group key exchange. Lecture Notes in Computer Science, 2729:110–125, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camenisch:2003:PVE

[140] Jan Camenisch and Victor Shoup. Practical verifiable encryption and decryption of discrete logarithms. Lecture Notes in Computer Science, 2729: 126–144, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishai:2003:EOT

[141] Yuval Ishai, Joe Kilian, Kobbi Nissim, and Erez Petrank. Extending oblivious transfers efficiently. Lecture Notes in Computer Science, 2729:145–161, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armknecht:2003:AAC

[142] Frederik Armknecht and Matthias Krause. Algebraic attacks on combiners with memory. Lecture Notes in Computer Science, 2729:162–175, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courtois:2003:FAA

[143] Nicolas T. Courtois. Fast algebraic attacks on stream ciphers with linear feedback. Lecture Notes in Computer Science, 2729:176–194, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biryukov:2003:CS

[144] Alex Biryukov, Christophe De Cannière, and Gustaf Dellkrantz. Cryptanalysis of Safer++. Lecture Notes in Computer Science, 2729:195–211, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheon:2003:PTA

[145] Jung Hee Cheon and Byungheup Jun. A polynomial time algorithm for the braid Diffie–Hellman conjugacy problem. Lecture Notes in Computer Science, 2729:212–225, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howgrave-Graham:2003:IDF

[146] Nick Howgrave-Graham, Phong Q. Nguyen, David Pointcheval, John Proos, Joseph H. Silverman, Ari Singer, and William Whyte. The impact of decryption failures on the security of NTRU encryption. Lecture Notes in Computer Science, 2729:226–246, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damgaard:2003:UCE

[147] Ivan Damgård and Jesper Buus Nielsen.

Universally composable efficient multiparty computation from threshold homomorphic encryption. Lecture Notes in Computer Science, 2729:247–264, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canetti:2003:UCJ

[148] Ran Canetti and Tal Rabin. Universal composition with joint state. Lecture Notes in Computer Science, 2729: 265–281, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Micciancio:2003:SZK

[149] Daniele Micciancio and Salil P. Vadhan. Statistical zero-knowledge proofs with efficient provers: Lattice problems and more. Lecture Notes in Computer Science, 2729:282–298, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barak:2003:DC

[150] Boaz Barak, Shien Jin Ong, and Salil Vadhan. Derandomization in cryptography. Lecture Notes in Computer Science, 2729:299–315, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pass:2003:DCR

[151] Rafael Pass. On deniability in the common reference string and random oracle model. Lecture Notes in Computer Science, 2729:316–337, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:2003:PPO

[152] Qi Cheng. Primality proving via one round in ECPP and one iteration in AKS. Lecture Notes in Computer Science, 2729:338–348, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rubin:2003:TBC

[153] Karl Rubin and Alice Silverberg. Torusbased cryptography. Lecture Notes in Computer Science, 2729:349–365, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Komano:2003:EUP

[154] Yuichi Komano and Kazuo Ohta. Efficient universal padding techniques for multiplicative trapdoor one-way permutation. Lecture Notes in Computer Science, 2729:366–382, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyen:2003:MIB

[155] Xavier Boyen. Multipurpose identitybased signcryption: A Swiss Army knife for identity-based cryptography. Lecture Notes in Computer Science, 2729: 383–399, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krawczy:2003:SSM

[156] Hugo Krawczy. SIGMA: The “SIGnand-MAc” approach to authenticated Diffie–Hellman and its use in the IKE protocols. Lecture Notes in Computer Science, 2729:400–425, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dwork:2003:MBF

[157] Cynthia Dwork, Andrew Goldberg, and Moni Naor. On memory-bound functions for fighting spam. Lecture Notes in Computer Science, 2729:426–444, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchbinder:2003:LUB

[158] Niv Buchbinder and Erez Petrank.

Lower and upper bounds on obtaining history independence. Lecture Notes in Computer Science, 2729:445–462, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishai:2003:PCS

[159] Yuval Ishai, Amit Sahai, and David Wagner. Private circuits: Securing hardware against probing attacks. Lecture Notes in Computer Science, 2729: 463–481, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halevi:2003:TEM

[160] Shai Halevi and Phillip Rogaway. A tweakable enciphering mode. Lecture Notes in Computer Science, 2729: 482–499, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cary:2003:MAC

[161] Matthew Cary and Ramarathnam Venkatesan. A message authentication code based on unimodular matrix groups. Lecture Notes in Computer Science, 2729:500–512, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patarin:2003:LRR

[162] Jacques Patarin. Luby–Rackoff: 7 rounds are enough for security. Lecture Notes in Computer Science, 2729: 513–529, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horvitz:2003:WKA

[163] Omer Horvitz and Virgil Gligor. Weak key authenticity and the computational completeness of formal encryption. Lecture Notes in Computer Science, 2729: 530–547, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzog:2003:PAK

[164] Jonathan Herzog, Moses Liskov, and Silvio Micali. Plaintext awareness via key registration. Lecture Notes in Computer Science, 2729:548–564, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canetti:2003:RCC

[165] Ran Canetti, Hugo Krawczyk, and Jesper B. Nielsen. Relaxing chosenciphertext security. Lecture Notes in Computer Science, 2729:565–582, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canvel:2003:PIS

[166] Brice Canvel, Alain Hiltgen, Serge Vaudenay, and Martin Vuagnoux. Password interception in a SSL/TLS channel. Lecture Notes in Computer Science, 2729:583–599, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barkan:2003:ICO

[167] Elad Barkan, Eli Biham, and Nathan Keller. Instant ciphertext-only cryptanalysis of GSM encrypted communication. Lecture Notes in Computer Science, 2729:600–616, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oechslin:2003:MFC

[168] Philippe Oechslin. Making a faster cryptanalytic time-memory trade-off. Lecture Notes in Computer Science, 2729:617–630, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stajano:2003:SCU

[169] Frank Stajano. The security challenges of ubiquitous computing. Lecture Notes in Computer Science, 2779:1, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:2003:MCA

[170] Dakshi Agrawal, Josyula R. Rao, and Pankaj Rohatgi. Multi-channel attacks. Lecture Notes in Computer Science, 2779:2–16, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karlof:2003:HMM

[171] Chris Karlof and David Wagner. Hidden Markov model cryptanalysis. Lecture Notes in Computer Science, 2779:17–34, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ors:2003:PAA

[172] Sıddıka Berna Örs, Elisabeth Oswald, and Bart Preneel. Power-analysis attacks on an FPGA — first experimental results. Lecture Notes in Computer Science, 2779:35–50, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geiselmann:2003:HSS

[173] Willi Geiselmann and Rainer Steinwandt. Hardware to solve sparse systems of linear equations over GF(2). Lecture Notes in Computer Science, 2779:51–61, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsunoo:2003:CIC

[174] Yukiyasu Tsunoo, Teruo Saito, Tomoyasu Suzaki, Maki Shigeri, and Hiroshi Miyauchi. Cryptanalysis of DES implemented on computers with cache. Lecture Notes in Computer Science, 2779:62–76, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piret:2003:DFA

[175] Gilles Piret and Jean-Jacques Quisquater. A differential fault attack technique against SPN structures, with application to the AES and KHAZAD. Lecture Notes in Computer Science, 2779:77–88, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coron:2003:NAS

[176] Jean-Sébastien Coron and Alexei Tchulkine. A new algorithm for switching from arithmetic to Boolean masking. Lecture Notes in Computer Science, 2779:89–97, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:2003:DNP

[177] Jovan D. Golić. DeKaRT: A new paradigm for key-dependent reversible circuits. Lecture Notes in Computer Science, 2779:98–112, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karri:2003:PBC

[178] Ramesh Karri, Grigori Kuznetsov, and Michael Goessel. Parity-based concurrent error detection of substitutionpermutation network block ciphers. Lecture Notes in Computer Science, 2779: 113–124, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiri:2003:SEA

[179] Kris Tiri and Ingrid Verbauwhede. Securing encryption algorithms against DPA at the logic level: Next generation Smart Card technology. Lecture Notes in Computer Science, 2779:125–136, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fournier:2003:SEA

[180] Jacques J. A. Fournier, Simon Moore, Huiyun Li, Robert Mullins, and George Taylor. Security evaluation of asynchronous circuits. Lecture Notes in Computer Science, 2779:137–151, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epstein:2003:DIT

[181] Michael Epstein, Laszlo Hars, Raymond Krasinski, Martin Rosner, and Hao Zheng. Design and implementation of a true random number generator based on digital circuit artifacts. Lecture Notes in Computer Science, 2779:152–165, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barak:2003:TRN

[182] Boaz Barak, Ronen Shaltiel, and Eran Tromer. True random number generators secure in a changing environment. Lecture Notes in Computer Science, 2779:166–180, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dichtl:2003:HPO

[183] Markus Dichtl. How to predict the output of a hardware random number generator. Lecture Notes in Computer Science, 2779:181–188, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reyhani-Masoleh:2003:LCB

[184] Arash Reyhani-Masoleh and M. Anwar Hasan. On low complexity bit parallel polynomial basis multipliers. Lecture Notes in Computer Science, 2779: 189–202, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dhem:2003:EMR

[185] Jean-François Dhem. Efficient modular reduction algorithm in and its application to “left to right” modular multiplication in. Lecture Notes in Computer Science, 2779:203–213, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chevallier-Mames:2003:FDS

[186] Benoît Chevallier-Mames, Marc Joye, and Pascal Paillierinst. Faster doublesize modular multiplication from Euclidean multipliers. Lecture Notes in Computer Science, 2779:214–227, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwon:2003:EEC

[187] Soonhak Kwon, Chang Hoon Kim, and Chun Pyo Hong. Efficient exponentiation for a class of finite fields GF(2^{n}) determined by Gauss periods. Lecture Notes in Computer Science, 2779: 228–242, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joye:2003:GFA

[188] Marc Joye and Pascal Paillier. GCDfree algorithms for computing modular inverses. Lecture Notes in Computer Science, 2779:243–253, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fouque:2003:AUR

[189] Pierre-Alain Fouque, Gwenaëlle Martinet, and Guillaume Poupard. Attacking unbalanced RSA–CRT using SPA. Lecture Notes in Computer Science, 2779:254–268, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fouque:2003:DAW

[190] Pierre-Alain Fouque and Frederic Valette. The doubling attack — why upwards is better than downwards. Lecture Notes in Computer Science, 2779: 269–280, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smart:2003:AGR

[191] Nigel P. Smart. An analysis of Goubin’s refined power analysis attack. Lecture Notes in Computer Science, 2779: 281–290, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cathalo:2003:NTT

[192] Julien Cathalo, François Koeune, and Jean-Jacques Quisquater. A new type of timing attack: Application to GPS. Lecture Notes in Computer Science, 2779: 291–303, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satoh:2003:UHA

[193] Akashi Satoh and Sumio Morioka. Unified hardware architecture for 128-bit block ciphers AES and Camellia. Lecture Notes in Computer Science, 2779: 304–318, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chodowiec:2003:VCF

[194] Paweł Chodowiec and Kris Gaj. Very compact FPGA implementation of the AES algorithm. Lecture Notes in Computer Science, 2779:319–333, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Standaert:2003:EIR

[195] Francois-Xavier Standaert, Gael Rouvroy, Jean-Jacques Quisquater, and Jean-Didier Legat. Efficient implementation of Rijndael encryption in reconfigurable hardware: Improvements and design tradeoffs. Lecture Notes in Computer Science, 2779:334–350, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelzl:2003:HCC

[196] Jan Pelzl, Thomas Wollinger, Jorge Guajardo, and Christof Paar. Hyperelliptic curve cryptosystems: Closing the performance gap to elliptic curves. Lecture Notes in Computer Science, 2779: 351–365, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avanzi:2003:CAD

[197] Roberto M. Avanzi. Countermeasures against differential power analysis for hyperelliptic curve cryptosystems. Lecture Notes in Computer Science, 2779: 366–381, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itoh:2003:PCA

[198] Kouichi Itoh, Tetsuya Izu, and Masahiko Takenaka. A practical countermeasure against address-bit differential power analysis. Lecture Notes in Computer Science, 2779:382–396, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okeya:2003:MFC

[199] Katsuyuki Okeya and Tsuyoshi Takagi. A more flexible countermeasure against side channel attacks using window method. Lecture Notes in Computer Science, 2779:397–410, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clulow:2003:SP

[200] Jolyon Clulow. On the security of PKCS #11. Lecture Notes in Computer Science, 2779:411–425, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klima:2003:ARB

[201] Vlastimil Klíma, Ondrej Pokorný, and Tomá š Rosa. Attacking RSA-based sessions in SSL/TLS. Lecture Notes in Computer Science, 2779:426–440, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).